# X-SHIELD: Regularization for eXplainable Artificial Intelligence

### v2

@@ 2.3 Towards the Integration of XAI via Regularization Methods
This generalization manages to incorporate into the regularization term constraints that we want the model to have regarding the specific dataset, without allowing these constraints to bias the final prediction. In (Weber et al., 2023), different ways of improving models based on explanations are theoretically defined and classified depending on the part of the machine learning algorithm is used to improve. Within this classification of this work, the addition of regularization terms to the loss functions is categorized as a Loss Augmentation approach to improve explainability.
@@ 3.1 T-SHIELD regularization family
Formally, for , a certain hide transformation can be described as , where if is one of the features not hided and if not, where is a neutral value that can be chosen.
@@
The definition of T-SHIELD regularization family has no theoretical restrictions except that the model must be differentiable. Besides, T-SHIELD regularization can be applied to any type of input, output, model, and task. It should also be noted that the use of T-SHIELD regularization can be classified as a Loss Augmentation method on the classification of XAI-based augmentation methods proposed in (Weber et al., 2023).
@@ 3.2 X-SHIELD: selection of the transformation T to enhance model explainability
In this section, we define the X-SHIELD regularization with the selection of a certain XAI-inspired transformation. The main goal of the proposed X-SHIELD transformation is to introduce knowledge by the action of hiding specific features. We use XAI techniques to hide the most appropriate features for this purpose.
@@
The choice of the features to be hided is based on the estimation of its importance. Formally, and as described in (Sevillano-García et al., 2023), the absolute importance of a feature in feature importance explanations can be estimated as , where is the vector of the component of the importance matrix, and operator is the module operator of a vector. This factor is then the value of the differentiation of the model output with respect to the feature . For the sake of lower training time and simplicity, the approach chosen to determine the importance of each feature is the aforementioned approach. This approach is also implemented in different artificial intelligence libraries since it is used to the gradient descendant optimization algorithm.
With the previous definition of absolute importance of each feature , let be the transformation fixed to the value . The parameter is fixed in and the features are sorted by increasing absolute importance, thus making the first least-important features to be hidden.
@@ 3.3 Technical aspects for T-SHIELD regularizations
Once the concept of features in images is defined, we propose a method of hiding these features from the model we are working with. This task is non-trivial due to the current lack of models specifically designed for this purpose. A commonly quick and simple way used for this task is to change the region of the selected superpixels to a neutral color of choice, either black, white or the mean of the image or superpixel (Amparore et al., 2021). We implement in the proposal the method where we change the superpixel region to the mean color of the whole image.
@@ 4.1 Benchmark selection
The datasets selected as benchmarks are shown in Table 2 with a small analytical description. This selection is due partly to the fact that certain of these datasets have been used as benchmarks also in (Yan et al., 2022) and (Sevillano-García et al., 2023) for explainability tasks on image datasets. We added Flowers and Oxford IIIT Pet to study datasets with fewer examples. There is variety in terms of image size, dataset size, number of classes, and black and white vs. color images which allows us to draw general conclusions from the study.
@@
| Oxford IIIT Pet | 37 | 3680 | 3669 | Yes | (Parkhi et al., 2012) |
@@ 4.2 Models and optimization techniques
As base models, we use Efficientnet (Tan and Le, 2019) and Efficientnet V2 (Tan and Le, 2021) architectures since they are the state-of-the-art image classification datasets in their larger versions (Efficientnet B7 or Efficientnet V2-Medium). However, since the aim of this work is not to optimize the explanations, we decided to use lighter and faster models in terms of training time. We choose Efficientnet B2 and Efficientnet V2 Small architectures to check whether the study is not dependent on a particular model. We use Adam optimization technique (Kingma and Ba, 2014) with L2 regularization for a of 32.
Each dataset has Test and Training sets separated, determined a priori by the specific dataset. Each Training set has been separated into a 90% training set and a 10% validation set. The training set is used to update the weights of the model on the training step. The validation set is used to validate every epoch of the loss function and maintain the best-results model weights for testing. We run each experiment for 80 epochs.
@@
To test the performance of X-SHIELD and R-SHIELD, we use the same features as in LIME: we divide in 64 squares of the same size each image. To study the influence of on the X-SHIELD and R-SHIELD regularizations, we set and for Section 5.1. As a result of this section, we conclude that a lower value ends up in better results. For this reason, we then choose and for the rest of the experiments.
@@ 5.1 Impact of Nº of features hided ( parameter)
We show in Figures 3 and 4, the evolution graphs of loss, accuracy and regularization in the training and validation sets of X-SHIELD and R-SHIELD regularizations. We observe in both experiments that the lower the value, the better the model’s evolution graph in terms of loss and accuracy. This is expected since the parameter controls how many features are hidden. Therefore, the lower the value, the lower the difference between the original and the modified example is. This behavior leads to an earlier convergence. However, we notice that both regularizations decrease the convergence speed. The regularization terms adds complexity with the extra term to the loss function which explains this difficulty to learn.
The regularization term is higher when using X-SHIELD instead of R-SHIELD, as we may note in Figures 3(e) and 4(e), where both graphics shows similar behavior with different scales. X-SHIELD regularization maximum value is around and R-SHIELD maximum value is around . We must highlight that X-SHIELD regularization term is deterministic while R-SHIELD regularization term is sthocastic. As a consecuence, R-SHIELD gives the same relevance sthocastically to all features, while X-SHIELD do not give relevance to the features that the model interpret as not important.
@@ 5.2.1 Introduction of the X-SHIELD regularization analysis
Table 3 shows the comparison of a basic training and the use of X-SHIELD in terms of accuracy and loss metrics. We observe that, in terms of accuracy, the best configuration is achieved by X-SHIELD in 11 out of 12 experiments, with basic training achieving the best score in 1 case. However, in terms of the loss metric, the basic training does not achieve the best score in any case. This might indicate that using X-SHIELD regularization makes de model generalize better.
@@
| V2 S | X-SHIELD | 15 | 93.35% | 0.0075 |
@@
As a final visualization of the results, in Figure 6 we show the 3D barchart plots with the ’s values and the different accuracy error rates over different datasets.
@@ 5.2.2 Selection of the transformation T in performance enhancement
| V2 S | X | 15 | 93.35% | 0.0075 |
To check if these experiments have significant differences, we show in Figure 7 the Bayesian Signed-Rank tests. As both p-value are lower than 95%, the differences between X-SHIELD and R-SHIELD are not significant, even if X-SHIELDS have a higher p-value. Then, we conclude that X-SHIELD and R-SHIELD has similar performance. However, we still need to evaluate the explainability perspective since it is the main improvement introduced by X-SHIELD regularization.

### v3

@@ 2.3 Towards the Integration of XAI via Regularization Methods
This generalization manages to incorporate into the regularization term constraints that we want the model to have regarding the specific dataset, without allowing these constraints to bias the final prediction. In (Weber et al., 2023), different ways of improving models based on explanations are theoretically defined and classified depending on the part of the machine learning algorithm is used to improve. Within this classification of this work, the addition of regularization terms to the loss functions is categorized as a Loss Augmentation approach to improve explainability. There are two different approaches, depending on the need of human-explanations.
On the one hand, there are works that require explanations in order to ensure that a decision has been made for the right reasons. These approaches have the disadvantage of relying on human-labelled explanations. The obtention of these explanations is difficult and time-consuming because it requires human labeling besides the classification labeling (Ross et al., 2017).
On the other hand, other works propose to improve explanations without prior labeled explanations. These approaches propose to improve the characteristics of the explanations, without knowing whether these explanations are the ones that a human would provide, such as Saliency Guided (Ismail et al., 2021) and SaliencyMix (Uddin et al., 2020). These approaches do not require prior labeling. However, these approaches use specific explanations of the assigned class. Moreover, the analysis of those methods are measured exclusively with the module of gradients as a measure of quality for explainability. To make sure of the improvement in explainability, it is mandatory to use quantitative descriptive measures of the resulting explanations.
@@ 3.1 T-SHIELD regularization family
Formally, for , a certain hide transformation can be described as , where if is one of the features not hided and if not, where is a neutral value that can be chosen. In Section 3.3, we detail the complexities of choosing and our approach to choose it as a hiding technique.
@@
The definition of T-SHIELD regularization family has no theoretical restrictions except that the model must be differentiable. Besides, T-SHIELD regularization can be applied to any type of input, output, model, and task. It should also be noted that the use of T-SHIELD regularization can be classified as a Loss Augmentation method on the classification of Loss augmentation methods proposed in (Weber et al., 2023). Also, unlike Saliency Guided (Ismail et al., 2021), T-SHIELD use a commutative loss function between the input and the modified input. This property makes our proposal to balance the influence of the input and the modified input instead of prioritizing that one of the outputs resembles the other.
@@ 3.2 X-SHIELD: selection of the transformation T to enhance model explainability
In this section, we define the X-SHIELD regularization with the selection of a certain XAI-inspired transformation. The main goal of the proposed X-SHIELD transformation is to introduce knowledge by the action of hiding specific features. We use XAI techniques to hide the most appropriate features for this purpose, categorizing this proposal as a technique for improving the model with Loss Augmentation without human explanation labeling.
@@
The choice of the features to be hided is based on the estimation of its importance. Formally, and as described in (Sevillano-García et al., 2023), the absolute importance of a feature in feature importance explanations can be estimated as where is the vector of the component of the importance matrix, and operator is the module operator of a vector. This factor is then the value of the differentiation of the model output with respect to the feature . For the sake of lower training time and simplicity, the approach chosen to determine the importance of each feature is the aforementioned approach. This approach is also implemented in different artificial intelligence libraries since it is used to the gradient descendant optimization algorithm.
With the previous definition of absolute importance of each feature , let be the transformation fixed to the value . The parameter is fixed in and the features are sorted by increasing absolute importance, thus making the first least-important features to be hidden. Our regularization function can be formalized as follows:
Note that this approach uses a general explanation of the model decision. This generalization allows it to be generalized as a regularization function in any task.
@@ 3.3 Technical aspects for T-SHIELD regularizations
Once the concept of features in images is defined, we propose a method of hiding these features from the model we are working with. Following the formulation of 3.1, the hiding technique can be understood as the by which to replace particular parts of the image. This task is non-trivial due to the current lack of models specifically designed for this purpose. A commonly quick and simple way used for this task is to change the region of the selected superpixels to a neutral color of choice, either black, white or the mean of the image or superpixel (Amparore et al., 2021). We implement in the proposal the method where we change the superpixel region to the mean color of the whole image.
@@ 4.1 Benchmark selection
The datasets selected as benchmarks are shown in Table 2 with a small analytical description. This selection is due partly to the fact that certain of these datasets have been used as benchmarks also in (Yan et al., 2022) and (Sevillano-García et al., 2023) for explainability tasks on image datasets. We added Flowers and Oxford IIIT Pet to study datasets with fewer examples. We also added the large-scale dataset of Imagenet 1K to test the effectiveness of this regularization on datasets with this amount of data. There is variety in terms of image size, dataset size, number of classes, and black and white vs. color images which allows us to draw general conclusions from the study.
@@
Oxford IIIT Pet | 37 | 3680 | 3669 | Yes | (Parkhi et al., 2012) | |
Imagenet 1k | 1000 | 1.281.167 | 100.000 | Yes | (Deng et al., 2009) |
@@ 4.2 Models and optimization techniques
As base models, we use Efficientnet (Tan and Le, 2019) and Efficientnet V2 (Tan and Le, 2021) architectures since they are the state-of-the-art image classification datasets in their larger versions (Efficientnet B7 or Efficientnet V2-Medium). However, since the aim of this work is not to optimize the explanations, we decided to use lighter and faster models in terms of training time. We choose Efficientnet B2 and Efficientnet V2 Small architectures to check whether the study is not dependent on a particular model. We use Adam optimization technique (Kingma and Ba, 2014) with L2 regularization for a of 32. To perform the experiments, we use a Tesla A100 GPU with 8 cores.
Each dataset has Test and Training sets separated, determined a priori by the specific dataset. Each Training set has been separated into a 90% training set and a 10% validation set. The training set is used to update the weights of the model on the training step. The validation set is used to validate every epoch of the loss function and maintain the best-results model weights for testing. We run each experiment for 80 epochs. We do not use this configuration of epoch in the Imagenet 1k training because of the amount of data it has and the computational cost: one epoch on Imagenet 1k consumes around 1 day. In this case, we consider an epoch to be updating weights of 5000 batches to decrease the computation time. The validation epoch has not been modified.
@@
To test the performance of X-SHIELD and R-SHIELD, we use the same features as in LIME: we divide in 64 squares of the same size each image. To study the influence of on the X-SHIELD and R-SHIELD regularizations, we set and for Section 5.1. As a result of this section, we conclude that a lower value ends up in better results. For this reason, we then choose and for the rest of the experiments. Due to the computational time discussed in this section, for the experimentation with Imagenet 1k we have only considered for X-SHIELD and R-SHIELD. This specific case is discussed separately.
@@ 5.1 Impact of Nº of features hided ( parameter)
We show in Figures 3(f) and 4(f), the evolution graphs of loss, accuracy and regularization in the training and validation sets of X-SHIELD and R-SHIELD regularizations. We observe in both experiments that the lower the value, the better the model’s evolution graph in terms of loss and accuracy. This is expected since the parameter controls how many features are hidden. Therefore, the lower the value, the lower the difference between the original and the modified example is. This behavior leads to an earlier convergence. However, we notice that both regularizations decrease the convergence speed. The regularization terms adds complexity with the extra term to the loss function which explains this difficulty to learn.
The regularization term is higher when using X-SHIELD instead of R-SHIELD, as we may note in Figures 3(e) and 4(e), where both graphics shows similar behavior with different scales. X-SHIELD regularization maximum value is around and R-SHIELD maximum value is around . We must highlight that X-SHIELD regularization term is deterministic while R-SHIELD regularization term is sthocastic. As a consecuence, R-SHIELD gives the same relevance stochastically to all features, while X-SHIELD do not give relevance to the features that the model interpret as not important.
@@ 5.2.1 Introduction of the X-SHIELD regularization analysis
Table 3 shows the comparison of a basic training and the use of X-SHIELD in terms of accuracy and loss metrics. We observe that, in terms of accuracy, the best configuration is achieved by X-SHIELD in 13 out of 14 experiments, with basic training achieving the best score in 1 case. However, in terms of the loss metric, the basic training does not achieve the best score in any case. This might indicate that using X-SHIELD regularization makes de model generalize better.
@@
V2 S | X-SHIELD | 15 | 93.35% | 0.0075 | |
Basic training | 0 | 75.66% | 0.0314 | ||
Imagenet 1K | B2 | X-SHIELD | 3* | 76.09% | 0.0295 |
Basic training | 0 | 78.87% | 0.0279 | ||
V2 S | X-SHIELD | 3* | 79.89% | 0.0249 |
@@
As a final visualization of the results, in Figure 6 we show the 3D barchart plots with the ’s values and the different accuracy error rates over different datasets but for Imagenet 1K due to cost and time restrictions.
@@
Finally, and without leaving aside the Imagenet 1k dataset, we note in Table 3 that X-SHIELD also performs better than the base training on this dataset with .
@@ 5.2.2 Selection of the transformation T in performance enhancement
V2 S | X | 15 | 93.35% | 0.0075 | |
R | 0 | 75.82% | 0.0298 | ||
Imagenet 1K | B2 | X | 3* | 75.86% | 0.0295 |
R | 0 | 79.89% | 0.0248 | ||
V2 S | X | 3* | 79.89% | 0.0249 |
To check if these experiments have significant differences, we show in Figure 7 the Bayesian Signed-Rank tests. As both p-value are lower than 95%, the differences between X-SHIELD and R-SHIELD are not significant, even if X-SHIELDS have a higher p-value of 0.895 against 0.103. Then, we conclude that X-SHIELD and R-SHIELD has similar performance. However, we still need to evaluate the explainability perspective since it is the main improvement introduced by X-SHIELD regularization.